Hi this is a question in my summer review packet. I have forgotton how to figure it out if you could help that would be great.
The radius of a circle is 15 centimeters. Find the arc length of a sector with an arc measure of 120 degrees.

Answer by checkley71(8403)

circumferance=2*3.14*15 or 94.2 cn. 120/360=1/3 of the circumferance of the circle thus 94.2*1/3=94.2/3=31.4 cn